The car with mass m = 2t passes at a speed of 30 km / h through a bridge bent upwards with radius R = 12 m. Calculate the compressive force exerted by the car on the bridge at the highest point.


1
Expert's answer
2021-01-19T07:10:17-0500
"F=mg+m\frac{v^2}{r}\\
F=(2000)(9.8)+2000\frac{30^2}{12(3.6)^2}\\F=31000\ N=31\ kN"